OpenStudy (anonymous):

In right triangle ABC, if A=15 degrees, c=37 and angle c is the right angle. Solve the triangle. Round to the nearest tenth, if necessaary.

jigglypuff314 (jigglypuff314):

well to get angle B know that there are a total of 180 degrees in a triangle angle C = 90 (definition of right angle) angle A = 15 so angle B = 180 - 90 - 15 =

OpenStudy (anonymous):

i got 75

OpenStudy (anonymous):

but what about the 37

jigglypuff314 (jigglypuff314):

yep :) so the over all problem looks like 

jigglypuff314 (jigglypuff314):

then you can use cos to find side b cos(15) = sideb/37 so sideb = cos(15) times 37 =

jigglypuff314 (jigglypuff314):

and you can find side a by sin(15) = side a / 37 so side a = sin(15) times 37 =

OpenStudy (anonymous):

9.576304668811

jigglypuff314 (jigglypuff314):

yup a = "9.576304668811" then b = 37(cos(15)) =

OpenStudy (anonymous):

11.979532650705